5.152g of Na2CO3 (s) is added to 75.0 mL of water at 22.00 C. When the dissolution is complete, the final temperature of the solution is measured at 25.80 C. The final volume of the solution is 75.8 mL, and the density of water at 25.21 C is 0.99780 g/ml. Calculate the vapor pressure ( in mm Hg) of the solution.
1
Expert's answer
2019-09-17T07:42:21-0400

The vapor pressure of water at 25.0 °C is 23.76 mm Hg. We need the mole fraction of Na2CO3 in water in order to use Raoult's Law.

n(Na2CO3)=5.152g/106g/mol=0.0486mol

m(H2O)=0.9978g/ml*75ml=74.835g

n(H2O)=74.835g/18g/mol=4.1575mol

Total moles is 4.2061.

Mole fraction=0.0486mol/4.1575mol=0.01169

∆P=mole fraction*P0=0.01169*23.76mm Hg= 0.278 mm Hg.

Vapor pressure of the solution=23.76-0.278=23.483 mm Hg.